# On anomalous dissipation induced by transport noise

### v2

@@ 1.3.4. Further comparison with the literature
As commented in Subsection 1.3.2, the main novelty consists of an improvement of the scaling limit (1.15). Let us note that such an improvement is also interesting on its own. Indeed, the latter can be used also to strengthen other results on the topic. For instance, arguing as in Section 4, one can check that the main results of [32] can be extended also to the case of $L^{2}(\mathbb{T}^{d})$ being critical for the corresponding SPDE (here, we use criticality in the sense of [7, Section 3] and are implicitly assuming that [32, (H4)] holds with $L^{2}$ replaced by $H^{r}$ with $r>0$). For instance, this covers the case of the Keller-Segel model system in four dimensions, cf., [32, Subsection 2.1].
@@
We are now in a position to give a comparison between our results and the one of [49, Section 3] and [73]. Our results are in spirit similar to the one in [73] which, roughly speaking, ensure that anomalous dissipation can happen if the noise is chosen according to the viscosity and/or diffusivity (cf., the comments below [73, Theorem 1.1]). On the one hand, in our results, due to the application of the scaling limit in (1.14)-(1.15), we cannot provide an explicit dependence on $\mathrm{supp}\,\theta^{\gamma}$ on $\gamma$ while the one in [73] are rather explicit. On the other hand, in our results, we have a sharp control of how the intensity of the noise affects anomalous dissipation and our methods are robust enough for applications to nonlinear SPDEs, e.g., 2D NSEs in vorticity formulation leading to anomalous dissipation of enstrophy as conjectured in the KB theory. From a technical point of view, the approaches are completely different and thus a comparison is not possible.
@@
In [49], the authors showed total dissipations for passive scalars and for 2D and (even) 3D weak solutions to the NSEs. Here, total dissipation means that all the initial energy $\|\varrho_{0}\|_{L^{2}}>0$ is absorbed by the energy dissipation rates, i.e., $\lim_{t\uparrow 1}\|\varrho^{\gamma}(t)\|_{L^{2}}=0$ a.s. for all $\gamma>0$. Our results share some similarities with the one of [49, Section 3], where the authors considered transport noise. For instance, to the best of our knowledge, they were the first to use the constructions in [32, 44] in the context of anomalous dissipation. However, there are several differences between our results and the one in [49, Section 3], and in particular in the mechanism for creating anomalous dissipation. Firstly, instead of assuming a viscosity/diffusivity dependent $\theta$ in either (1.1) or (1.3), they considered a piecewise constant in time noise coefficient $\theta(t)=(\theta_{k}(t))_{k\in{\mathbb{Z}}^{d}_{0}}$ such that $\|\theta(t)\|_{\ell^{2}}\to\infty$ as $t\uparrow 1$, see [49, Subsection 2.1]. In particular, following the argument in Subsection 1.2, extending the identity (1.10) to the case of piecewise constant in time $\theta$, one sees that their noise has ‘infinite energy’ as $t\uparrow 1$. This is not the case in our construction where, due to the requirement $\theta^{\nu}\in\mathcal{S}_{\ell^{2}}^{0}$ for all $\nu>0$, we obtain finite energy of the noise as $\nu\downarrow 0$ and uniformly for $t\in[0,1]$. However, our method cannot deal with 3D NSEs in contrast to [49] (see also comments below Theorem 1.1). Secondly, from a technical point of view, the proofs are completely different. Indeed, in [49, Section 3] the authors rely on a careful study of the mixing properties of the transport noise and subsequently upgraded them to an anomalous dissipation result. This approach in particular requires $\sup_{t\in[0,1)}\|\theta(t)\|_{\ell^{2}}=\infty$. In our manuscript, instead, we employ the above-discussed refinement of the scaling limit arguments of [44, 32] via Meyers’ estimates. To conclude, it would be interesting to see if the arguments used in [49, Section 4] to derive anomalous dissipation by advection via randomly forced NSEs can benefit from the use of Meyers’ estimates. However, this goes beyond the scope of the current manuscript.
@@ 2.1. Notation
Next, we collect the notation for function spaces. We write $L^{p}(S,\mu;X)$ for the Bochner space of strongly measurable, $p$-integrable $X$-valued functions for a measure space $(S,\mu)$ and a Banach space $X$ as defined in [51, Section 1.2b]. If $X={\mathbb{R}}$, we write $L^{p}(S,\mu)$ and if it is clear which measure we refer to we also leave out $\mu$. Finally, $\mathbf{1}_{A}$ denotes the indicator function of $A\subseteq S$. Bessel-potential spaces are indicated as usual by $H^{s,q}(\mathbb{T}^{d})$ where $s\in{\mathbb{R}}$ and $q\in(1,\infty)$. We also use the standard shorth-hand notation $H^{s,2}(\mathbb{T}^{d})$ for $H^{s}(\mathbb{T}^{d})$. We sometimes also employ Besov spaces $B^{s}_{q,p}(\mathbb{T}^{d})$ which can be defined as the real interpolation space $(H^{-k,q}(\mathbb{T}^{d}),H^{k,q}(\mathbb{T}^{d}))_{(s+k)/2k,p}$ for $s\in{\mathbb{R}}$, ${\mathbb{N}}\ni k>|s|$ and $1<q,p<\infty$. The reader is referred to [16, 51] for details on interpolation and to [74, Section 3.5.4] for details on function spaces over $\mathbb{T}^{d}$. Finally, we let $\mathcal{A}(\mathbb{T}^{d};{\mathbb{R}}^{k})\stackrel{{\scriptstyle{\rm def}}}{{=}}(\mathcal{A}(\mathbb{T}^{d}))^{k}$ and $\mathcal{A}(\cdot)\stackrel{{\scriptstyle{\rm def}}}{{=}}\mathcal{A}(\mathbb{T}^{d};\cdot)$ for $\mathcal{A}\in\{L^{q},H^{s,q},B^{s}_{q,p}\}$ and $k\in{\mathbb{N}}$.
2.2. Structure of the noise
Here, we describe the quantities $\theta^{\nu}$, $\theta^{\gamma}$, $\sigma_{k,\alpha}$, $W^{k,\alpha}$ and $W^{k}$ appearing in the stochastic perturbation of (1.1) and (1.3), while $c_{d}\stackrel{{\scriptstyle{\rm def}}}{{=}}d/(d-1)$. In this subsection, we follow [32, 37]. In the following, we only describe the stochastic perturbation in (1.3), for the case of NSEs (1.1) it is enough to take $d=2$ in the following construction and omit the index $\alpha$. Moreover, since $\gamma>0$ is fixed here, we simply write $\theta$ instead of $\theta^{\gamma}$.
To begin, set ${\mathbb{Z}}_{0}^{d}\stackrel{{\scriptstyle{\rm def}}}{{=}}{\mathbb{Z}}^{d}\setminus\{0\}$. Throghout the manuscript $\theta=(\theta_{k})_{k\in{\mathbb{Z}}^{d}_{0}}\in\ell^{2}({\mathbb{Z}}^{d}_{0})$ is normalized and radially symmetric, i.e.
(2.1) $\|\theta\|_{\ell^{2}({\mathbb{Z}}^{d}_{0})}=1\quad\text{ and }\quad\theta_{j}=\theta_{k}\ \text{ for all $j,k\in{\mathbb{Z}}_{0}^{d}$ \ such that }|j|=|k|.$
Next, we define the family of vector fields $(\sigma_{k,\alpha})_{k,\alpha}$. Here and in the following, we use the shorthand subscript ‘$k,\alpha$’ to indicate $k\in{\mathbb{Z}}^{d}_{0},\,\alpha\in\{1,\dots,d-1\}$. Let ${\mathbb{Z}}_{+}^{d}$ and ${\mathbb{Z}}_{-}^{d}$ be a partition of ${\mathbb{Z}}_{0}^{d}$ such that $-{\mathbb{Z}}_{+}^{d}={\mathbb{Z}}_{-}^{d}$. For any $k\in{\mathbb{Z}}_{+}^{d}$, let $\{a_{k,\alpha}\}_{\alpha\in\{1,\dots,d-1\}}$ be a complete orthonormal basis of the hyperplane $k^{\bot}=\{k^{\prime}\in{\mathbb{R}}^{d}\,:\,k\cdot k^{\prime}=0\}$, and set $a_{k,\alpha}\stackrel{{\scriptstyle{\rm def}}}{{=}}a_{-k,\alpha}$ for $k\in{\mathbb{Z}}^{d}_{-}$. Finally, we let
@@ 2.2. Structure of the noise
As in [37, Section 2.3] or [32, Remark 1.1], by (2.1) and the definition of the vector fields $\sigma_{k,\alpha}$, at least formally, one has
@@
where we recall that $c_{d}=d/(d-1)$. To see the above, one uses that $\nabla\cdot\sigma_{k,\alpha}=0$, (2.1) and the elementary identity (cf. [37, eq. (2.3)] or [32, eq. (3.2)])
@@
Let us remark that the stochastic integration on the RHS(2.3) is understood in the Itô–sense. In the paper we will always understand the Stratonovich noise on the LHS(2.3) as the RHS(2.3), namely an Itô noise plus a diffusion term. However, note that the diffusion term $\mu\Delta\varrho$ does not provide any additional diffusion, as in energy estimates it balances the Itô correction coming from the Itô-noise, cf., (1.13).
@@ 2.3. Solution concept and well-posedness
$\mathcal{W}_{\ell^{2}}(f)=\sum_{k,\alpha}\int_{{\mathbb{R}}_{+}}f_{k,\alpha}(t)\,\mathrm{d}W^{k,\alpha}_{t}\ \ \ \text{ for }\ \ \ f=(f_{k,\alpha})_{k,\alpha}\in L^{2}({\mathbb{R}}_{+};\ell^{2}).$
@@
We begin with defining solutions to the passive scalar equation (1.3). Below we use the reformulation of the Stratonovich noise in (1.3).
@@
Fix $\gamma>0$, $\theta^{\gamma}\in\ell^{2}$ and $\varrho_{0}\in L^{2}(\mathbb{T}^{d})$. Let $\tau^{\gamma}:\Omega\to[0,\infty]$ and $\varrho^{\gamma}:[0,\tau^{\gamma})\times\Omega\to\mathbb{H}^{1}(\mathbb{T}^{d})$ be a stopping time and a progressive measurable process, respectively.
  • We say that $(\varrho^{\gamma},\tau^{\gamma})$ is a local solution to (1.3) if the following are satisfied:
    • $\varrho^{\gamma}\in L^{2}_{{\rm loc}}([0,\tau^{\gamma});H^{1}(\mathbb{T}^{d}))\cap C([0,\tau^{\gamma});L^{2}(\mathbb{T}^{d}))$ a.s.;
    • a.s. for all $t\in[0,\tau^{\gamma})$ it holds that
@@
      $+$ $\,\sqrt{c_{d}\mu}\int_{0}^{t}\mathbf{1}_{[0,\tau^{\gamma})}\Big{(}(\theta^{\gamma}_{k}\,\sigma_{k,\alpha}\cdot\nabla)\varrho^{\gamma}(s)\Big{)}_{k,\alpha}\,\mathrm{d}\mathcal{W}_{\ell^{2}}.$
  • A local solution $(\varrho^{\gamma},\tau^{\gamma})$ to (1.3) is a said to be a unique local solution to (1.3) if for any local solution $(\chi^{\gamma},\lambda^{\gamma})$ we have $\lambda^{\gamma}\leq\tau^{\gamma}$ a.s. and $\chi^{\gamma}=\varrho^{\gamma}$ a.e. on $[0,\lambda^{\gamma})\times\Omega$.
  • A unique local solution $(\varrho^{\gamma},\tau^{\gamma})$ to (1.3) is said to be a unique global solution to (1.3) if $\tau^{\gamma}=\infty$ a.s.
@@
Existence and uniqueness of global solutions to (1.3) in the sense of Definition 2.1 is standard, see e.g., [59, Chapter 4]. In case $\theta^{\gamma}_{k}$ decays sufficiently fast as $|k|\to\infty$ (e.g., #$\#\{k\,:\,\theta_{k}^{\gamma}\neq 0\}<\infty$), then the unique global solution $\varrho^{\gamma}$ instantaneously improves its regularity in time and space, cf.  [6, Theorems 2.7 and 4.2].
@@
Fix $\nu>0$, $\theta^{\nu}\in\ell^{2}$ and $\zeta_{0}\in L^{2}(\mathbb{T}^{2})$. Let $\tau^{\nu}:\Omega\to[0,\infty]$ and $\zeta^{\nu}:[0,\infty)\times\Omega\to H^{1}(\mathbb{T}^{2})$ be a stopping time and a progressive measurable process, respectively.
  • We say that $(\zeta^{\nu},\tau^{\nu})$ is a local solution to (1.1) if the following are satisfied:
    • $\zeta^{\nu}\in L^{2}_{{\rm loc}}([0,\infty);H^{1}(\mathbb{T}^{2}))\cap C([0,\infty);L^{2}(\mathbb{T}^{2}))$ a.s.;
    • ${\mathcal{K}}\,\zeta^{\nu}\zeta^{\nu}\in L^{2}_{{\rm loc}}([0,\infty);L^{2}(\mathbb{T}^{2};{\mathbb{R}}^{2}))$ a.s.;
    • a.s. for all $t\in{\mathbb{R}}_{+}$ it holds that
@@
      $+\sqrt{2\mu}\int_{0}^{t}\mathbf{1}_{[0,\tau^{\nu})}\Big{(}\theta^{\nu}_{k}\,(\sigma_{k}\cdot\nabla)\zeta^{\nu}(s)\Big{)}_{k}\,\mathrm{d}\mathcal{W}_{\ell^{2}}.$
  • A local solution $(\zeta^{\nu},\tau^{\nu})$ to (1.1) is said to be a unique local solution to (1.1) if for any local solution $(\xi^{\nu},\lambda^{\nu})$ we have $\lambda^{\nu}\leq\tau^{\nu}$ a.s. and $\xi^{\nu}=\zeta^{\nu}$ a.e. on $[0,\lambda^{\nu})\times\Omega$.
  • A unique local solution $(\zeta^{\nu},\tau^{\nu})$ to (1.1) is said to be a unique global solution to (1.1) if $\tau^{\nu}=\infty$ a.s.
As above, the integrals in Definition 2.2 are well-defined and, in case of global solutions, we only write $\zeta^{\nu}$ instead of $(\zeta^{\nu},\tau^{\nu})$ if no confusion seems likely.
Existence and uniqueness of global solutions to (1.1) is standard, see e.g., either [7, Theorem 3.4] or [40]. Instantaneous regularity results in case $\theta^{\nu}_{k}$ decays sufficiently fast as $|k|\to\infty$ can be proven as in [9, Theorems 2.7 and 2.12].

### v3

@@ 1.3.4. Further comparison with the literature
As commented in Subsection 1.3.2, the main novelty consists of an improvement of the scaling limit (1.17). Let us note that such an improvement is also interesting on its own. Indeed, the latter can also be used to strengthen other results on the topic. For instance, arguing as in Section 4, one can check that the main results of [36] can be extended also to the case of $L^{2}(\mathbb{T}^{d})$ being critical for the corresponding SPDE (here, we use criticality in the sense of [8, Section 3] and are implicitly assuming that [36, (H4)] holds with $L^{2}$ replaced by $H^{r}$ with $r>0$). For instance, this covers the case of the Keller-Segel model system in four dimensions, cf. [36, Subsection 2.1].
@@
We are now in a position to give a comparison between our results and those of [55, Section 3] and [80]. Our results are, in spirit, similar to the ones in [80] which, roughly speaking, ensure that anomalous dissipation can happen if the noise is chosen according to the viscosity and/or diffusivity (cf. the comments below [80, Theorem 1.1]). On the one hand, in our results, due to the application of the scaling limit in (1.16)-(1.17), we cannot provide an explicit dependence on $\mathrm{supp}\,\theta^{\gamma}$ on $\gamma$ while those in [80] are rather explicit. On the other hand, in our results, we have a sharp control of how the intensity of the noise affects anomalous dissipation, and our methods are robust enough for applications to nonlinear SPDEs, e.g., 2D NSEs in vorticity formulation leading to anomalous dissipation of enstrophy as conjectured in the KB theory. From a technical point of view, the approaches are completely different and thus a comparison is not possible.
@@
In [55], the authors showed total dissipation for passive scalars and 2D and (even) 3D weak solutions to the NSEs. Here, total dissipation means that all the initial energy $\|\varrho_{0}\|_{L^{2}}>0$ is absorbed by the energy dissipation rates, i.e., $\lim_{t\uparrow 1}\|\varrho^{\gamma}(t)\|_{L^{2}}=0$ a.s. for all $\gamma>0$. Our results share some similarities with those of [55, Section 3], where the authors considered transport noise. For instance, to the best of our knowledge, they were the first to use the constructions in [36, 49] in the context of anomalous dissipation. However, there are several differences between our results and those in [55, Section 3], and in particular in the mechanism for creating anomalous dissipation. Firstly, instead of assuming a viscosity/diffusivity dependent $\theta$ in either (1.1) or (1.3), they considered a piecewise constant in time noise coefficient $\theta(t)=(\theta_{k}(t))_{k\in{\mathbb{Z}}^{d}_{0}}$ such that $\|\theta(t)\|_{\ell^{2}}\to\infty$ as $t\uparrow 1$, see [55, Subsection 2.1]. In particular, following the argument in Subsection 1.2, extending the identity (1.12) to the case of piecewise constant in time $\theta$, one sees that their noise has ‘infinite energy’ as $t\uparrow 1$. This is not the case in our construction where, due to the requirement $\theta^{\nu}\in\mathcal{S}_{\ell^{2}}^{0}$ for all $\nu>0$, we obtain finite energy of the noise as $\nu\downarrow 0$ and uniformly for $t\in[0,1]$. However, our method cannot deal with 3D NSEs in contrast to [55] (see also comments below Theorem 1.1). Secondly, from a technical point of view, the proofs are completely different. Indeed, in [55, Section 3], the authors rely on a careful study of the mixing properties of the transport noise and subsequently upgrade them to an anomalous dissipation result. This approach in particular requires $\sup_{t\in[0,1)}\|\theta(t)\|_{\ell^{2}}=\infty$. In our manuscript, instead, we employ the above-discussed refinement of the scaling limit arguments of [49, 36] via Meyers’ estimates. To conclude, it would be interesting to see if the arguments used in [55, Section 4] to derive anomalous dissipation by advection via randomly forced NSEs can benefit from the use of Meyers’ estimates. However, this goes beyond the scope of the current manuscript.
@@ 2.1. Notation
Below, $(\Omega,\mathscr{A},(\mathscr{F}_{t})_{t\geq 0},\mathbf{P})$ denotes a filtered probability space carrying a sequence of independent standard Brownian motions which changes depending on the SPDE under consideration. We write $\mathbf{E}$ for the expectation on $(\Omega,\mathscr{A},\mathbf{P})$ and $\mathscr{P}$ for the progressive $\sigma$-field. A process $\phi:[0,\infty)\times\Omega\to X$ is progressively measurable if $\phi|_{[0,t]\times\Omega}$ is $\mathscr{B}([0,t])\otimes\mathscr{F}_{t}$ measurable for all $t\geq 0$, where $\mathscr{B}$ is the Borel $\sigma$-algebra on $[0,t]$ and $X$ a Banach space. Moreover, a stopping time $\tau$ is a measurable map $\tau:\Omega\to[0,\infty]$ such that $\{\tau\leq t\}\in\mathscr{F}_{t}$ for all $t\geq 0$. For a stopping time $\tau$, we define the stochastic interval $0,\tau)\times\Omega$ as
$0,\tau)\times\Omega\stackrel{{\scriptstyle{\rm def}}}{{=}}\{(t,\omega)\in[0,\infty)\times\Omega\,:\,\,0\leq t<\tau(\omega)\}.$
A similar definition is employed for $0,\tau]\times\Omega$, $(0,\tau]\times\Omega$ etc. Finally, a stochastic process $\phi:[0,\tau)\times\Omega\to X$ is progressively measurable if $\mathbf{1}_{[0,\tau)\times\Omega}\,\phi$ is progressively measurable where $\mathbf{1}_{[0,\tau)\times\Omega}$ stands for the extension by zero outside the set $0,\tau)\times\Omega$, and $\tau$ is a stopping time.
Next, we collect the notation for function spaces. We write $L^{p}(S,\mu;X)$ for the Bochner space of strongly measurable, $p$-integrable $X$-valued functions for a measure space $(S,\mu)$ and a Banach space $X$ as defined in [57, Section 1.2b]. If $X={\mathbb{R}}$, we write $L^{p}(S,\mu)$, and if it is clear which measure we refer to, we also leave out $\mu$. Finally, $\mathbf{1}_{A}$ denotes the indicator function of $A\subseteq S$. Bessel-potential spaces are indicated as usual by $H^{s,q}(\mathbb{T}^{d})$ where $s\in{\mathbb{R}}$ and $q\in(1,\infty)$. We also use the standard shorth-hand notation $H^{s,2}(\mathbb{T}^{d})$ for $H^{s}(\mathbb{T}^{d})$. We sometimes also employ Besov spaces $B^{s}_{q,p}(\mathbb{T}^{d})$ which can be defined as the real interpolation space $(H^{-k,q}(\mathbb{T}^{d}),H^{k,q}(\mathbb{T}^{d}))_{(s+k)/2k,p}$ for $s\in{\mathbb{R}}$, ${\mathbb{N}}\ni k>|s|$ and $1<q,p<\infty$. The reader is referred to [18, 57] for details on interpolation and to [81, Section 3.5.4] for details on function spaces over $\mathbb{T}^{d}$. Finally, we let $\mathcal{A}(\mathbb{T}^{d};{\mathbb{R}}^{k})\stackrel{{\scriptstyle{\rm def}}}{{=}}(\mathcal{A}(\mathbb{T}^{d}))^{k}$ and $\mathcal{A}(\cdot)\stackrel{{\scriptstyle{\rm def}}}{{=}}\mathcal{A}(\mathbb{T}^{d};\cdot)$ for $\mathcal{A}\in\{L^{q},H^{s,q},B^{s}_{q,p}\}$ and $k\in{\mathbb{N}}$.
2.2. Structure of the noise
Here, we describe the quantities $\theta^{\nu}$, $\theta^{\gamma}$, $\sigma_{k,\alpha}$, $W^{k,\alpha}$ and $W^{k}$ appearing in the stochastic perturbation of (1.1) and (1.3), while $c_{d}\stackrel{{\scriptstyle{\rm def}}}{{=}}d/(d-1)$. In this subsection, we follow [36, 42]. In the following, we only describe the stochastic perturbation in (1.3), for the case of NSEs (1.1), it is enough to take $d=2$ in the following construction and omit the index $\alpha$. Moreover, since $\gamma>0$ is fixed here, we simply write $\theta$ instead of $\theta^{\gamma}$.
To begin, set ${\mathbb{Z}}_{0}^{d}\stackrel{{\scriptstyle{\rm def}}}{{=}}{\mathbb{Z}}^{d}\setminus\{0\}$. Throughout the manuscript $\theta=(\theta_{k})_{k\in{\mathbb{Z}}^{d}_{0}}\in\ell^{2}({\mathbb{Z}}^{d}_{0})$ is normalized and radially symmetric, i.e.
@@ 2.2. Structure of the noise
Next, we define the family of vector fields $(\sigma_{k,\alpha})_{k,\alpha}$. Here and in the following, we use the shorthand subscript $k,\alpha$’ to indicate $k\in{\mathbb{Z}}^{d}_{0},\,\alpha\in\{1,\dots,d-1\}$. Let ${\mathbb{Z}}_{+}^{d}$ and ${\mathbb{Z}}_{-}^{d}$ be a partition of ${\mathbb{Z}}_{0}^{d}$ such that ${\mathbb{Z}}_{+}^{d}={\mathbb{Z}}_{-}^{d}$. For any $k\in{\mathbb{Z}}_{+}^{d}$, let $\{a_{k,\alpha}\}_{\alpha\in\{1,\dots,d-1\}}$ be a complete orthonormal basis of the hyperplane $k^{\bot}=\{k^{\prime}\in{\mathbb{R}}^{d}\,:\,k\cdot k^{\prime}=0\}$, and set $a_{k,\alpha}\stackrel{{\scriptstyle{\rm def}}}{{=}}a_{-k,\alpha}$ for $k\in{\mathbb{Z}}^{d}_{-}$. Finally, we let
@@
As in [42, Section 2.3] or [36, Remark 1.1], by (2.1) and the definition of the vector fields $\sigma_{k,\alpha}$, at least formally, one has
@@
where we recall that $c_{d}=d/(d-1)$. To see the above, one uses that $\nabla\cdot\sigma_{k,\alpha}=0$, (2.1) and the elementary identity (cf. [42, eq. (2.3)] or [36, eq. (3.2)])
@@
Let us remark that the stochastic integration on the RHS(2.3) is understood in the Itô–sense. In the paper, we will always understand the Stratonovich noise on the LHS(2.3) as the RHS(2.3), namely an Itô noise plus a diffusion term. However, note that the diffusion term $\mu\Delta\varrho$ does not provide any additional diffusion, as in energy estimates it balances the Itô correction coming from the Itô-noise, cf. (1.15).
@@ 2.3. Solution concept and well-posedness
(2.5) $\mathcal{W}_{\ell^{2}}(f)=\sum_{k,\alpha}\int_{{\mathbb{R}}_{+}}f_{k,\alpha}(t)\,\mathrm{d}W^{k,\alpha}_{t}\ \ \ \text{ for }\ \ \ f=(f_{k,\alpha})_{k,\alpha}\in L^{2}({\mathbb{R}}_{+};\ell^{2}).$
@@
We begin by defining solutions to the passive scalar equation (1.3). Below, we use the reformulation of the Stratonovich noise in (1.3).
@@
Fix $\gamma>0$, $\theta^{\gamma}\in\ell^{2}$ and $\varrho_{0}\in L^{2}(\mathbb{T}^{d})$. Let $\tau^{\gamma}:\Omega\to[0,\infty]$ and $\varrho^{\gamma}:[0,\tau^{\gamma})\times\Omega\to H^{1}(\mathbb{T}^{d})$ be a stopping time and a progressive measurable process, respectively.
  • We say that $(\varrho^{\gamma},\tau^{\gamma})$ is a local solution to (1.3) if there exists a sequence of stopping times $(\tau^{\gamma}_{n})_{n\geq 1}$ such that $\tau^{\gamma}_{n}\leq\tau^{\gamma}$ a.s. for all $n\geq 1$, and $\lim_{n\to\infty}\tau^{\gamma}_{n}=\tau^{\gamma}$ a.s., for which the following conditions hold for all $n\geq 1$:
    • $\varrho^{\gamma}\in L^{2}(0,\tau_{n}^{\gamma};H^{1}(\mathbb{T}^{d}))\cap C([0,\tau_{n}^{\gamma}];L^{2}(\mathbb{T}^{d}))$ a.s.;
    • a.s. for all $t\in[0,\tau^{\gamma}_{n}]$ it holds that
@@
      $+$ $\,\sqrt{c_{d}\mu}\int_{0}^{t}\mathbf{1}_{[0,\tau^{\gamma}_{n}]}\Big((\theta^{\gamma}_{k}\,\sigma_{k,\alpha}\cdot\nabla)\varrho^{\gamma}(s)\Big)_{k,\alpha}\,\mathrm{d}\mathcal{W}_{\ell^{2}}.$
  • A local solution $(\varrho^{\gamma},\tau^{\gamma})$ to (1.3) is a said to be a unique maximal (local) solution to (1.3) if for any other local solution $(\chi^{\gamma},\lambda^{\gamma})$ we have $\lambda^{\gamma}\leq\tau^{\gamma}$ a.s. and $\chi^{\gamma}=\varrho^{\gamma}$ a.e. on $0,\lambda^{\gamma})\times\Omega$.
  • A unique maximal solution $(\varrho^{\gamma},\tau^{\gamma})$ to (1.3) is said to be a unique global solution to (1.3) if $\tau^{\gamma}=\infty$ a.s.
@@
Existence and uniqueness of global solutions to (1.3) in the sense of Definition 2.1 is standard, see e.g., [65, Chapter 4]. In case $\theta^{\gamma}_{k}$ decays sufficiently fast as $|k|\to\infty$ (e.g., $\{k\,:\,\theta_{k}^{\gamma}\neq 0\}<\infty$), then the unique global solution $\varrho^{\gamma}$ instantaneously improves its regularity in time and space, cf. [7, Theorems 2.7 and 4.2].
@@
Fix $\nu>0$, $\theta^{\nu}\in\ell^{2}$ and $\zeta_{0}\in L^{2}(\mathbb{T}^{2})$. Let $\tau^{\nu}:\Omega\to[0,\infty]$ and $\zeta^{\nu}:[0,\tau^{\nu})\times\Omega\to H^{1}(\mathbb{T}^{2})$ be a stopping time and a progressive measurable process, respectively.
  • We say that $(\zeta^{\nu},\tau^{\nu})$ is a local solution to (1.1) if there exists a sequence of stopping times $(\tau^{\nu}_{n})_{n\geq 1}$ such that $\tau^{\nu}_{n}\leq\tau^{\nu}$ a.s. for all $n\geq 1$, and $\lim_{n\to\infty}\tau^{\nu}_{n}=\tau^{\nu}$ a.s., for which the following conditions hold for all $n\geq 1$:
    • $\zeta^{\nu}\in L^{2}(0,\tau^{\nu}_{n};H^{1}(\mathbb{T}^{2}))\cap C([0,\tau^{\nu}_{n}];L^{2}(\mathbb{T}^{2}))$ a.s.;
    • ${\mathcal{K}}\,\zeta^{\nu}\zeta^{\nu}\in L^{2}(0,\tau^{\nu}_{n};L^{2}(\mathbb{T}^{2};{\mathbb{R}}^{2}))$ a.s.;
    • a.s. for all $t\in[0,\tau^{\nu}_{n}]$ it holds that
@@
      $+\sqrt{2\mu}\int_{0}^{t}\mathbf{1}_{[0,\tau^{\nu}_{n}]}\Big(\theta^{\nu}_{k}\,(\sigma_{k}\cdot\nabla)\zeta^{\nu}(s)\Big)_{k}\,\mathrm{d}\mathcal{W}_{\ell^{2}}.$
  • A local solution $(\zeta^{\nu},\tau^{\nu})$ to (1.1) is said to be a unique maximal (local) solution to (1.1) if for any other local solution $(\xi^{\nu},\lambda^{\nu})$ we have $\lambda^{\nu}\leq\tau^{\nu}$ a.s. and $\xi^{\nu}=\zeta^{\nu}$ a.e. on $0,\lambda^{\nu})\times\Omega$.
  • A unique maximal solution $(\zeta^{\nu},\tau^{\nu})$ to (1.1) is said to be a unique global solution to (1.1) if $\tau^{\nu}=\infty$ a.s.
As above, the integrals in Definition 2.2 are well-defined and, in the case of global solutions, we only write $\zeta^{\nu}$ instead of $(\zeta^{\nu},\tau^{\nu})$ if no confusion seems likely.
Existence and uniqueness of global solutions to (1.1) is standard, see e.g., [8, Theorem 3.4], [11, Theorem 7.10] or [45]. Instantaneous regularization results in case $\theta^{\nu}_{k}$ decays sufficiently fast as $|k|\to\infty$ can be proven as in [10, Theorems 2.7 and 2.12].